# Explainable AI: Comparative Analysis of Normal and Dilated ResNet Models for Fundus Disease Classification

### v1

@@ Abstract
This paper presents dilated Residual Network (ResNet) models for disease classification from retinal fundus images. Dilated convolution filters are used to replace normal convolution filters in the higher layers of the ResNet model (dilated ResNet) in order to improve the receptive field compared to the normal ResNet model for disease classification. This study introduces computer-assisted diagnostic tools that employ deep learning, enhanced with explainable AI techniques. These techniques aim to make the tool’s decision-making process transparent, thereby enabling medical professionals to understand and trust the AI’s diagnostic decision. They are particularly relevant in today’s healthcare landscape, where there is a growing demand for transparency in AI applications to ensure their reliability and ethical use. The dilated ResNet is used as a replacement for the normal ResNet to enhance the classification accuracy of retinal eye diseases and reduce the required computing time. The dataset used in this work is the Ocular Disease Intelligent Recognition (ODIR) dataset which is a structured ophthalmic database with eight classes covering most of the common retinal eye diseases. The evaluation metrics used in this work include precision, recall, accuracy, and F1 score. In this work, a comparative study has been made between normal ResNet models and dilated ResNet models on five variants namely ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152. The dilated ResNet model shows promising results as compared to normal ResNet with an average F1 score of 0.71, 0.70, 0.69, 0.67, and 0.70 respectively for the above respective variants in ODIR multiclass disease classification.
@@
journal: Biomedical Signal Processing and Control
[…]
[inst1]organization=Computer Science and Engineering,addressline=Indian Institute of Information Technology Design and Manufacturing Kancheepuram, city=Chennai, postcode=600127, state=Tamil Nadu, country=India
@@ 5.3 Results
Table 2: Results of ResNet-18 with and without dilation
@@
Table 3: Results of ResNet-34 with and without dilation
@@
Table 4: Results of ResNet-50 with and without dilation
@@
Table 5: Results of ResNet-101 with and without dilation
@@
Table 6: Results of ResNet-152 with and without dilation
@@ 5.4 Explainable AI Techniques
(a) Input image (Class: Glaucoma)
@@
(a) Input image (Class: Diabetes)
@@
(a) Input image (Class: Diabetes)
(b) Without dilation: Predicted - cataract
(c) With dilation: Predicted - diabetes
@@ 5.5 Analysis of Activation Maps of ResNet Models
Figure 14 shows the activation map obtained for the class cataract with and without dilation. Similarly, Figure 15 and Figure 21 shows the activation map obtained for the classes normal and abnormalities with and without dilation. It can be observed that the activation maps obtained with dilation correctly focus on the entire region for three classes (cataract, normal, and abnormalities), in contrast to those without dilation.
Figure 16 shows the activation map obtained for the class glaucoma with and without dilation. Similarly, Figure 19 and Figure 18 shows the activation map obtained for the classes myopia and degeneration with and without dilation. It is clear from the comparison with dilation focus on the optic disc and near optic disc (macula) correctly when compared to without dilation.
Figure 17 shows the activation map obtained for the class diabetic retinopathy with and without dilation. Similarly, Figure 20 shows the activation map obtained for the class hypertension with and without dilation. It is clear from the comparison with dilation focus on the with dilation focus on blood vessels correctly when compared to without dilation.

### v2

@@ Abstract
This paper presents dilated Residual Network (ResNet) models for disease classification from retinal fundus images. Dilated convolution filters are used to replace normal convolution filters in the higher layers of the ResNet model (dilated ResNet) in order to improve the receptive field compared to the normal ResNet model for disease classification. This study introduces computer-assisted diagnostic tools that employ deep learning, enhanced with explainable AI techniques. These techniques aim to make the tool’s decision-making process transparent, thereby enabling medical professionals to understand and trust the AI’s diagnostic decision. They are particularly relevant in today’s healthcare landscape, where there is a growing demand for transparency in AI applications to ensure their reliability and ethical use. The dilated ResNet is used as a replacement for the normal ResNet to enhance the classification accuracy of retinal eye diseases and reduce the required computing time. The dataset used in this work is the Ocular Disease Intelligent Recognition (ODIR) dataset which is a structured ophthalmic database with eight classes covering most of the common retinal eye diseases. The evaluation metrics used in this work include precision, recall, accuracy, and F1 score. In this work, a comparative study has been made between normal ResNet models and dilated ResNet models on five variants namely ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152. The dilated ResNet model shows promising results as compared to normal ResNet with an average F1 score of 0.71, 0.70, 0.69, 0.67, and 0.70 respectively for the above respective variants in ODIR multiclass disease classification..
@@ 5.3 Results
Table 2: Results of ResNet-18 with and without dilation
@@
Table 3: Results of ResNet-34 with and without dilation
@@
Table 4: Results of ResNet-50 with and without dilation
@@
Table 5: Results of ResNet-101 with and without dilation
@@
Table 6: Results of ResNet-152 with and without dilation
@@ 5.4 Explainable AI Techniques
(a) Input image (Glaucoma)
@@
(a) Input image (Diabetes)
@@
(a) Input image (Diabetes)
(b) With dilation: Predicted as diabetes
(c) Without dilation: predicted as cataract
@@ 5.5 Analysis of Activation Maps of ResNet Models
Figure 14 shows the activation maps obtained for the class cataract with and without dilation. Similarly, Figure 15 and Figure 21 shows the activation maps obtained for the classes normal and abnormalities with and without dilation. It can be observed that the activation maps obtained with dilation correctly focus on the entire region for three classes (cataract, normal, and abnormalities), in contrast to those without dilation.
Figure 16 shows the activation maps obtained for the class glaucoma with and without dilation. Similarly, Figure 19 and Figure 18 shows the activation maps obtained for the classes myopia and degeneration with and without dilation. It is clear from the comparison with dilation focus on the optic disc and near optic disc (macula) correctly when compared to without dilation.
Figure 17 shows the activation maps obtained for the class diabetic retinopathy with and without dilation. Similarly, Figure 20 shows the activation maps obtained for the class hypertension with and without dilation. It is clear from the comparison with dilation focus on the with dilation focus on blood vessels correctly when compared to without dilation.
@@ 6 Conclusions
7 Declarations
Conflict of Interest: The authors declare that they have no conflict of interest.
7.1 Data Availability
The ODIR dataset used in the study is publicly available.
It can be downloaded from
- https://odir2019.grand-challenge.org/
7.2 Authors’ Contributions
  1. 1.
    P.N. Karthikayan: Literature survey, Training and testing, Verification, Drafting
  2. 2.
    Yoga Sri Varshan V: Algorithm design, Conceptual design, Implementation, Analysis, Verification
  3. 3.
    Hitesh Gupta Kattamuri: Training and testing, Verification, Analysis
  4. 4.
    Umarani Jayaraman: Verification, Drafting, Supervision